f(x +h)-f(x) By determining f(x) = lim h h0 find f(5) for the given function. f(x) = 6x2 f(5)=(Simplify your answer.)
Answer #1

solution: By definition, f(2)= lim f(a+h)-f(x) 2-0 f(n)=6x² =) + (5)= ? f(n) = 672 we have {(x)= lim {6(2th) - 6x²} ho = lim